# The associative-poset point of view on right regular bands

### v1

@@ 1 Introduction
This paper is a continuation of our joint work with A. Petrovich [2], where we study right-regular bands (RRB) $(P,\cdot)$ using as a guiding concept the partial order which arise naturally from the stipulation
@@
We present two results on the relation between the class of RRBs and that of associative posets. The first one is motivated by the question on how to obtain a band from a given poset. To give some context, choosing an RRB structure for a given associative poset is not trivial, since this necessarily involves the Axiom of Choice as it was seen in [2, Thm. 6.2]. The answer we present here is to construct a left adjoint to the obvious forgetful functor that takes an RRB $(P,\cdot)$ to $(P,\leq)$ using (1). The construction of such a left adjoint is actually generalized to any class of relational structures $(X,R)$ defined from a variety using a finite conjunction of identities to define $R$. This is presented in Section 3, where we also discuss some other concrete cases of this construction (Section 3.1); notably, there is a corresponding adjunction between the category of RRBs and that of sets with an equivalence relation (Example 3.15).
@@ 2 Preliminaries
We remind the reader that a right regular band (RRB) is an idempotent semigroup satifying
@@
We state two simple lemmas whose proofs are left as an exercise for the interested reader, or can be checked at [2, Sect. 2].
@@ 3 An adjunction
The categories of right-regular bands and associative posets are canonically linked by a forgetful functor. An adjunction between them can be constructed explicitly by making use of the algebraic structure of RRBs and associative posets. Instead of only doing this, in this section we present a more general result, by constructing left adjoints for a broad family of forgetful functors that includes the one mentioned above.
@@ 3.1 Examples
In this example we show that an appropriate expansion $\mathcal{K}$ of the class of right-regular bands has $\mathcal{K}_{R}$ equal to the class of sets equipped with some equivalence relation.
Let $L:=\{\cdot,\theta\}$ and let $\mathcal{K}$ be the class satisfying all the right-regular band axioms together with the additional condition
@@
We have that $\mathcal{K}_{\text{alg}}$ is, once again, the class of right-regular bands. It can be proved that $\theta$ is always an equivalence relation.
@@
Let $L:=\{\cdot,\leq\}$ and $\mathcal{K}$ the class of commutative right-posemigroups. $\mathcal{K}_{\text{alg}}$ is the class of meet-semilattices (as algebras) and $\mathcal{K}_{R}$ is the class of meet-semilattices (as posets).
@@ 3.2 SAFT applicability
The variety of RRB’s is residually large.
@@
We will show that subdirectly irreducible right-regular bands of arbitrarily large cardinality exist. Fix an arbitrary set $X$ with two distinguished elements 0 and $1$. Let $A$ consist of the following functions from $X$ into itself:
@@
We conclude that there’s no cogenerating set for the variety of RRB’s, so the Special Adjoint Functor Theorem cannot be applied in this case.
4 Factor congruences of right-regular bands
@@ 5 Conclusion
We do not know if it is possible to obtain the $5$-cycle. We do have an example where one of its components is.
A final last remark is that in our seminal example involving RRBs (Corollary 3.13) we are considering plain poset homorphisms as the morphisms of the category $\mathsf{AP}$. We are very curious if there exists a more appropriate notion of morphism that preserves more of the structure of associative posets.

### v2

@@ 1 Introduction
This paper is a continuation of our joint work with A. Petrovich [2], where we study right regular bands (RRB) $(P,\cdot)$ using as a guiding concept the partial order which arise naturally from the stipulation
@@
We present two results on the relation between the class of RRBs and that of associative posets. The first one is motivated by the question on how to obtain a band from a given poset. To give some context, choosing an RRB structure for a given associative poset is not trivial, since this necessarily involves the Axiom of Choice as it was seen in [2, Thm. 6.2]. The answer we present here is to construct a left adjoint to the obvious forgetful functor that takes an RRB $(P,\cdot)$ to $(P,\leq)$ using (1). The construction of such a left adjoint is actually generalized to any class of relational structures $(X,R)$ obtained from a variety using a finite conjunction of identities to define $R$. This is presented in Section 3, where we also discuss some other concrete cases of this construction (Section 3.1); notably, there is a corresponding adjunction between the category of RRBs and that of sets with an equivalence relation (Example 3.15).
@@ 2 Preliminaries
We remind the reader that a right regular band (RRB) is an idempotent semigroup satisfying
@@
We state two simple lemmas whose proofs are left as exercises for the interested reader, or can be checked at [2, Sect. 2].
@@ 3 An adjunction
The categories of right regular bands and associative posets are canonically linked by a forgetful functor. An adjunction between them can be constructed explicitly by making use of the algebraic structure of RRBs and associative posets. Instead of only doing this, in this section we present a more general result, by constructing left adjoints for a broad family of forgetful functors that includes the one mentioned above.
@@ 3.1 Examples
In this example we show that an appropriate expansion $\mathcal{K}$ of the class of right regular bands has $\mathcal{K}_{R}$ equal to the class of sets equipped with some equivalence relation.
Let $L:=\{\cdot,\theta\}$ and let $\mathcal{K}$ be the class satisfying all the right regular band axioms together with the additional condition
@@
We have that $\mathcal{K}_{\text{alg}}$ is, once again, the class of right regular bands. It can be proved that $\theta$ is always an equivalence relation.
@@
Let $L:=\{\cdot,\leq\}$ and $\mathcal{K}$ the class of commutative right posemigroups. $\mathcal{K}_{\text{alg}}$ is the class of meet-semilattices (as algebras) and $\mathcal{K}_{R}$ is the class of meet-semilattices (as posets).
@@ 3.2 SAFT applicability
The variety of RRBs is residually large.
@@
We will show that subdirectly irreducible right regular bands of arbitrarily large cardinality exist. Fix an arbitrary set $X$ with two distinguished elements 0 and $1$. Let $A$ consist of the following functions from $X$ into itself:
@@
We conclude that there’s no cogenerating set for the variety of RRBs, so the Special Adjoint Functor Theorem cannot be applied in this case.
4 Factor congruences of right regular bands
@@ 5 Conclusion
It is not possible to obtain the $5$-cycle. We do not know if there is an example having it as one of its components.
A final last remark is that in our seminal example involving RRBs (Corollary 3.13) we are considering plain poset homomorphisms as the morphisms of the category $\mathsf{AP}$. We are very curious if there exists a more appropriate notion of morphism that preserves more of the structure of associative posets.